| #28 |
09 Feb 11:58 Log created
|
|
Start experiment
|
normal
|
|
|
| #29 |
09 Feb 12:19 Antenna VSWR
|
ryanque
|
Antennas
|
normal
|
|
|
| #30 |
10 Feb 07:12 Spark gap summary
|
A. Coronado
|
Spark Gaps
|
normal
|
|
|
| #31 |
10 Feb 07:15 Start of the experiment
|
I. Manay
|
Start experiment
|
normal
|
|
|
| #32 |
10 Feb 08:00 Transmitters fall twice
|
I. Manay
|
Transmitters
|
normal
|
|
|
| #33 |
10 Feb 12:30 Increase of the noise level
|
I. Manay
|
Plot
|
normal
|
|
|
| #34 |
10 Feb 17:40 Interference was observed
|
I. Manay
|
Plot
|
normal
|
|
|
| #35 |
10 Feb 17:45 Transmitters fall
|
I. Manay
|
Transmitters
|
normal
|
|
|
| #36 |
10 Feb 18:00 Imaging experiment is started
|
I. Manay
|
Acquisition system
|
normal
|
|
|
| #37 |
10 Feb 18:08 PC Ares is restarted
|
I. Manay
|
Reception
|
normal
|
|
|
| #39 |
10 Feb 19:59 Operation resume
|
I. Manay
|
General
|
normal
|
|
|
| #40 |
10 Feb 23:10 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #41 |
11 Feb 06:15 Spark gap summary
|
AVALDEZ
|
Spark Gaps
|
normal
|
|
|
| #42 |
11 Feb 07:00 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #43 |
11 Feb 07:00 Figures of Experiment
|
|
Plot
|
normal
|
|
|
| #44 |
11 Feb 06:15 Problems with channel E
|
AVALDEZ
|
Reception
|
normal
|
|
|
| #45 |
11 Feb 08:00 Operation Resume
|
AVALDEZ
|
General
|
normal
|
|
|
| #47 |
11 Feb 09:55 Spark gaps adjustment
|
A. Aguilar
|
Spark Gaps
|
normal
|
|
|
| #48 |
11 Feb 19:25 Transmitter power measurement
|
|
Transmitters
|
normal
|
|
|
| #49 |
11 Feb 11:11 Lost of data
|
A. Aguilar
|
Acquisition system
|
normal
|
|
|
| #50 |
11 Feb 12:20 Sun appears in data
|
A. Aguilar
|
Physical Phenomena
|
normal
|
|
|
| #51 |
11 Feb 14:14 Spark Gaps adjustment
|
A. Aguilar
|
Spark Gaps
|
normal
|
|
|
| #52 |
11 Feb 17:40 Imaging Experiment was started
|
A. Aguilar
|
Start experiment
|
normal
|
|
|
| #53 |
11 Feb 18:05 Spark Gaps Adjustment
|
A. Aguilar
|
Spark Gaps
|
normal
|
|
|
| #54 |
11 Feb 18:49 Transmitters go down
|
A. Aguilar
|
Transmitters
|
normal
|
|
|
| #55 |
11 Feb 19:52 Operation Resume
|
A. Aguilar
|
General
|
normal
|
|
|
| #56 |
11 Feb 23:00 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #57 |
11 Feb 20:31 Problem with PC ARES
|
|
Acquisition system
|
normal
|
|
|
| #58 |
11 Feb 20:32 Faraday Desincronization
|
|
Acquisition system
|
normal
|
|
|
| #59 |
11 Feb 20:32 Restart EW_Drift
|
|
Acquisition system
|
normal
|
|
|
| #60 |
11 Feb 20:40 Restart Faraday
|
|
Acquisition system
|
normal
|
|
|
| #61 |
11 Feb 23:00 Plots of rti and spectra
|
|
Process/Analysis Program
|
normal
|
|
|
| #62 |
12 Feb 01:40 OFF Imaging
|
|
Acquisition system
|
normal
|
|
|
| #63 |
12 Feb 03:30 Tx Off
|
|
Transmitters
|
normal
|
|
|
| #64 |
12 Feb 06:20 Moon Pass
|
|
General
|
normal
|
|
|
| #65 |
12 Feb 06:15 Spark gaps adjustment
|
|
Spark Gaps
|
normal
|
|
|
| #66 |
12 Feb 23:00 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #67 |
12 Feb 07:59 Operation Resume
|
|
General
|
normal
|
|
|
| #68 |
12 Feb 19:55 Transmitter power measurements
|
|
Transmitters
|
normal
|
|
|
| #79 |
13 Feb 19:20 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #81 |
13 Feb 23:59 Experiment plots
|
I. Manay
|
Plot
|
normal
|
|
|
| #82 |
14 Feb 01:30 Imaging experiment is finished
|
I. Manay
|
Acquisition system
|
normal
|
|
|
| #83 |
14 Feb 03:30 Transmitters fall
|
I. Manay
|
Transmitters
|
normal
|
|
|
| #84 |
14 Feb 06:30 Spark gaps summary
|
J. Lavado
|
Spark Gaps
|
normal
|
|
|
| #85 |
14 Feb 07:59 Operation resume
|
I. Manay
|
General
|
normal
|
|
|
| #86 |
14 Feb 19:10 Spark gaps summary
|
AVALDEZ
|
Spark Gaps
|
normal
|
|
|
| #87 |
14 Feb 11:50 Attenuation Channel E
|
AVALDEZ
|
Acquisition system
|
normal
|
|
|
| #88 |
14 Feb 19:25 Transmitter power measurement
|
|
Transmitters
|
normal
|
|
|
| #89 |
14 Feb 22:40 Restart Adquisition
|
P.Condor
|
Acquisition system
|
normal
|
|
|
| #91 |
15 Feb 07:05 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #92 |
14 Feb 19:45 Restart Digisonde
|
P. Condor
|
Other Instruments
|
normal
|
|
|
| #93 |
14 Feb 19:50 Faraday Desynchronization
|
P. Condor
|
Acquisition system
|
normal
|
|
|
| #94 |
14 Feb 20:10 Problem with Faraday and EW_Drifts
|
P. Condor
|
General
|
normal
|
|
|
| #95 |
14 Feb 20:43 Problem with Transmitter
|
P. Condor
|
Transmitters
|
normal
|
|
|
| #96 |
14 Feb 22:00 The problem was the pulse of EW_Drifts
|
P. Condor
|
Acquisition system
|
normal
|
|
|
| #97 |
14 Feb 22:40 Restart Acquisition
|
P. Condor
|
Acquisition system
|
normal
|
|
|
| #98 |
14 Feb 23:59 Plots of rti, noise and others
|
P. Condor
|
Process/Analysis Program
|
normal
|
|
|
| #99 |
15 Feb 02:07 Imaging OFF
|
P. Condor
|
Acquisition system
|
normal
|
|
|
| #100 |
14 Feb 23:58 Plot Imaging
|
|
Process/Analysis Program
|
normal
|
|
|
| #101 |
15 Feb 06:28 Problem with Tx
|
P. Condor
|
Transmitters
|
normal
|
|
|
| #102 |
15 Feb 06:40 Faraday Desynchronization
|
P. Condor
|
Acquisition system
|
normal
|
|
|
| #103 |
15 Feb 06:25 Spark gaps adjustment
|
A. Belleza
|
Spark Gaps
|
normal
|
|
|
| #104 |
15 Feb Operation Resume
|
P. Condor
|
General
|
normal
|
|
|
| #105 |
15 Feb 10:40 Spark Gaps Adjusment
|
|
Spark Gaps
|
normal
|
|
|
| #107 |
15 Feb 14:50 Spark Gaps Adjusment
|
|
Spark Gaps
|
normal
|
|
|
| #108 |
15 Feb 15:20 Change TR of channel A
|
|
Start experiment
|
normal
|
|
|
| #109 |
15 Feb Operation Resume
|
|
Start experiment
|
normal
|
|
|
| #110 |
15 Feb 19:00 Transmitter power measurements
|
|
Transmitters
|
normal
|
|
|
| #112 |
16 Feb 17:00 Operation Resumen
|
|
Start experiment
|
normal
|
|
|
| #113 |
18 Feb 15:00 Data uploaded to PUMA
|
I. Manay
|
General
|
normal
|
|
|